Operations Report – February 2nd

Crew 240 Operations Report 02-02-2022

SOL: 2

Name of person filing report: François Vinet

Non-nominal systems: NA

Notes on non-nominal systems: NA

ROVERS

Spirit rover used: no
Hours: NA
Beginning charge: NA
Ending charge: NA
Currently charging: (Commander’s note) Was unused, and we’re leaving under Atila’s management from now on.

Opportunity rover used: no
Hours: NA
Beginning charge: NA
Ending charge: NA
Currently charging: (Commander’s note) Was unused, and we’re leaving under Atila’s management from now on.

Curiosity rover used: yes
Hours: 165.6
Beginning charge: 100%
Ending charge: 97%
Currently charging: (Commander’s note) Was plugged in after EVA, we’re leaving it under Atila’s management from now on.

Perseverance rover used: yes
Hours: 226.6
Beginning charge: 100%
Ending charge: 94%
Currently charging: (Commander’s note) Was plugged in after EVA, we’re leaving it under Atila’s management from now on.

General notes and comments: NA

Summary of Hab operations:

WATER USE: 19 gallons

Water (static tank): 416 gallons
Water (loft tank): 35 gallons
Water Meter: 0155373.1 units
Static to Loft Pump used: yes
Static tank pipe heater (on or off): on
Static tank heater (On or off): on

Toilet tank emptied: no

Summary of internet: NA

Summary of suits and radios: The problem with suit #7 is fixed.

Summary of GreenHab operations:

WATER USE: 10 gallons at 8 am and 6 gallons at 6 pm

Heater: on

Supplemental light: on

Harvest: 52g of cherry tomatoes

Questions: NA

Summary of ScienceDome operations:

Dual split: Off

SOC: 24% at 6.50 pm (it will be checked every hour during the afternoon for the rest of the mission)

Summary of RAM operations: Suit #7 has been repaired (a connector at the switch level was slightly unscrewed)

Summary of any observatory issues: NA

Summary of health and safety issues: NA

Questions, concerns and requests to Mission Support: one burnable trash is waiting for pickup in the rear airlock. Thank you!